OpenStudy (anonymous):

Which value is the solution for the equation cotx/2=-1? A. 3pi/2 B. 3pi/4 C. 5pi/4 D. 7pi/4

OpenStudy (anonymous):

cot(x/2) or cot(x)/2?

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

First solve \[x/2=cot^{-1}(-1)\]. Then you have a normal equation for which you can solve x. Don't forget that there are multiple possible answers for arccot. It's of the form (...+k*pi with k element of Z)

OpenStudy (jdoe0001):

$$ \bf \large { cot\pmatrix{\frac{x}{2}} = -1\\ cot^{-1}\pmatrix{cot\pmatrix{\frac{x}{2}}} = cot^{-1}(-1)\\ \frac{x}{2} = \frac{3 \pi}{4}, \frac{7 \pi}{4}\\ } $$

hartnn (hartnn):

or you can plug in choices one by one and see which one satisfies your equation...

OpenStudy (anonymous):

Nvm totally wrong :p

OpenStudy (anonymous):

I can only pick one....

hartnn (hartnn):

lets try 1st one, cot (x/2) = cot (3pi/4) = ... ?

OpenStudy (jdoe0001):

$$ \frac{x}{2} = \frac{3 \pi}{4}, \frac{7 \pi}{4}\\ \frac{x}{2} = \frac{3 \pi}{4} \implies x = \frac{3 \pi}{2}\\ \frac{x}{2}=\frac{7 \pi}{4} \implies x = \frac{14 \pi}{4} $$

OpenStudy (anonymous):

I think its A

hartnn (hartnn):

correct :) because cot (3pi/4) = -1 ...

OpenStudy (jdoe0001):

yes, the other choice will be \(\cfrac{7\pi}{2}\) and is not in the choices

OpenStudy (anonymous):

A is correct!